OpenStudy (anonymous):

Which number would you add to one and nine-tenths to give an approximate sum of 6? three and one-third five and two-elevenths four and one-twelfth two and seven-eighths

hero (hero):

\[1\frac{9}{10} + x = 6\]\[\frac{19}{10} + x = 6\]\[x = 6 - \frac{19}{10}\]\[x = \frac{60}{10} - \frac{19}{10} \]\[x = \frac{41}{10}\]\[x = 4\frac{1}{10}\]

hero (hero):

Unfortunately, it wants an approximate fraction, not an exact one.